# GraphGhost: Tracing Structures Behind Large Language Models

### v1

@@ 3.3 Case study on GraphGhost
In the subgraph, we observe that some nodes are activated consistently across datasets, while others contribute only to specific domains or even individual samples. For example, the period symbol “.” at layer 22 triggers strong logit activations across all four datasets, and the token “So” frequently appears across different reasoning types. Domain-specific reasoning tokens also emerge within the same domain. For instance, the action “Adding” and arithmetic expressions such as “6+” appear in both GSM8K and MAWPS. In contrast, certain nodes are unique to individual samples—for example, $\mathrm{B_{18}}$ activates only in choice-option reasoning within the ARC dataset.
3.4 Analysis on GraphGhost
In Figure 4, we also observe two typical actions associated with the node $_{22}.$ The first is semantic understanding, exemplified by the “6+” operation. The second is logic shifting, where the activation leads to the token “So” in the subsequent prediction. We further analyze these two actions from a statistical perspective.
Semantic understanding
In Section 2.3, we discussed how edge information is merged within the inner layers. Building on this, we analyze the token-merging behavior of neuron nodes by examining their average in-degree across layers in LLMs’ settings. The statistical results, summarized in the Appendix D and illustrated in Figure 5, show the top tokens across different datasets and models. We observe that the tokens with the highest in-degree are domain-sensitive. For example, in the ProntoQA dataset, the model is more sensitive to the period token than to the comma. This suggests that LLMs activate specific neurons in response to domain-specific data to capture semantics. Across models, however, the activated neurons are generally consistent: tokens such as <space>, period, and “the" consistently appear among the top-ranked, and this alignment persists even after finetuning.
@@ 3.4 Analysis on GraphGhost
Figure 5: The top merging nodes across the models and datasets.
@@
We further analyze the merging behavior across layers by computing the in-degree ratio. Let $d_{i}^{(\ell)}$ denote the in-degree of neuron node $(t_{i},\ell)$, where $\ell$ is the layer index and $i$ indexes the neuron within that layer. The in-degree ratio of layer $\ell$ is then defined as $R_{\ell}=\frac{\sum_{i=1}^{N_{\ell}}d_{i}^{(\ell)}}{\sum_{k=1}^{L}\sum_{i=1}^{N_{k}}d_{i}^{(k)}},$ where $N_{\ell}$ is the number of neurons in layer $\ell$ and $L$ is the total number of layers. As shown in Figure 6, the in-degree ratio distribution is imbalanced across both models and tokens. Although Qwen models and the Llama model perform semantic merging primarily in the bottom and top layers, the exact layers differ between them. Similarly, token merging patterns are not always consistent. Most tokens are merged in shallow layers, whereas high in-degree tokens such as period and “the" tend to merge at higher layers.
@@
Logic shifting.
The reasoning structure is segmented by special tokens such as So, Step, and Wait Muennighoff et al. (2025); Bogdan et al. (2025). We assume that activated neuron nodes in the middle layers trigger these special token expressions at the logit layers. At the same time, the neuron nodes themselves are activated by input tokens. To capture both effects, we treat logit-layer nodes as signals directed back to the neuron nodes by reversing their edge directions. We then apply the PageRank algorithm to identify the most influential nodes. We summarize the results in Figure 7 and Figure 8, and the full details can be found in Appendix D.
Figure 7: Top contributing tokens and layers in Qwen3-0.6B in various datasets.
Figure 8: Top contributing tokens and layers across models evaluated by MAWPS dataset.
@@
4 GraphGhost affects the reasoning process
GraphGhost paves a new way in which LLM reasoning processes can be analyzed and controlled. To evaluate this, we perform perturbation experiments at two complementary levels. (i) At the input level, we modify the contributing tokens identified in Section 4.1 to examine how altering key linguistic elements affects the model’s reasoning trajectory and output behaviors. (ii) At the neuron level, we intervene on the influential neuron nodes detected via in-degree analysis and PageRank in Section 4.2, assessing how structural disruptions within the reasoning pathways influence the model’s internal logic flow. Together, these two perturbation perspectives reveal how surface-level token changes and internal structural shifts jointly shape the reasoning dynamics captured by GraphGhost.
4.1 Impact of contribution tokens
GraphGhost in LLMs leverages contributing tokens to capture context and perform reasoning. In practice, we assess their importance by masking tokens, replacing each with a blank token. To quantify this effect, we evaluate two metrics: Sparsity and Fidelity. Sparsity measures the proportion of contributing tokens in a given question, while Fidelity evaluates whether the answer generated from only the contributing tokens remains consistent with the original answer. Formally, $\mathrm{Sparsity}=\frac{\#C}{\#T}$ and $\mathrm{Fidelity}_{T}=\mathds{1}(LLM(C),LLM(T))$, where $C$ denotes the set of contributing tokens, $T$ is the full set of tokens, and $\mathds{1}$ is an indicator function returning 1 if the two answers are identical and 0 otherwise. In practice, we use gemini-2.5-flash-lite to identify whether the two predictions are the same, with the prompt as: "Comparing two answers for a question: Answer1 {LLM(T^)LLM($\hat${T})}, Answer2 {LLM(T^)LLM($\hat${T})}. Do they have the same final results?" To compare the effectiveness of GraphGhost across models, we evaluate 100 samples that are solvable by Qwen3-0.6B, Qwen2.5-1.5B, and DS-Qwen-1.5B. The results are presented in Figure 9.
@@ 4.1 Impact of contribution tokens
Figure 10: Changing the semantic understanding. The characters in changing the language case is translating the original answer.
As shown in Figure 9, compared with random masking of the given question, the contributing tokens identified by GraphGhost successfully preserve the generated answer. This effect is particularly significant in mathematical reasoning tasks such as MAWPS and GSM8K, where even when half of the question tokens are removed, LLMs are still able to produce the correct prediction. In contrast, the performance gap between randomly masked and selected tokens is smaller in scientific reasoning tasks such as ARC dataset. This is because in ARC the answers often involve selecting the correct option, and in some cases LLMs rely on direct recognition of the options from the given context rather than explicit reasoning.
4.2 Modification on the graph structure
@@
Case study
First, as discussed in Section 3.4, the impacts include changes in semantic understanding and logic shifting. As shown in Figure 10, the semantic understanding is reflected in changing the language formulation and word expression. When the neuron node “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share the same semantic signal. As a result, the output appears in Chinese, directly translated from English. Furthermore, in the word-expression change case, deleting the neuron “.” at layer 17 causes the token “/” to replace “÷.” This suggests that another neuron is responsible for propagating the division signal, expressing it as “/.” These findings indicate that LLMs store semantics redundantly across multiple neuron nodes: when the most relevant neuron is muted, the model relies on alternative neurons that produce similar expressions. In Figure 11, we demonstrate how GraphGhost influences reasoning by altering key steps or driving the generation into loops. In the key-step change case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer. In the loop case, deleting the neuron node “the” at layer 27 causes the reasoning to degenerate into a repetition, producing “theShaper Information.” These cases suggest that even a single neuron in the middle layer can fundamentally change the outcome of the reasoning process.
Change the reasoning structure Change the key steps Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk. Sam placed 4.0 crayons and 8.0 scissors on the desk. How many crayons are now there in total? Let’s think step by step (ground truth: 17) Original answer: … - Sam placed 4.0 crayons on the desk. - Sam also placed 8.0 scissors on the desk. …. - Total crayons on the desk = 7.0 + 4.0 + 8.0 - Total crayons on the desk = 19.0 Mute the neuron node ’.’ at layer 25 Perturbed answer: … - Sam placed 4.0 crayons on the desk. - So, the total number of crayons on the desk is now 6.0 (initial crayons) + 4.0 (Sam’s crayons) = 10.0 crayons. … - So, the total number of crayons = 7.0 (drawer) + 10.0 (desk) = 17.0 crayons.   Failed into loop Question: Alyssa picked 25.0 limes and Mike picked 32.0 limes. Tom picked 12.0 plums. How many limes were picked in all? Let’s think step by step Original answer: Step 1: Identify the number of limes picked by Alyssa and Mike. Alyssa picked 25.0 limes. Mike picked 32.0 limes. … Mute the neuron node ’the’ at layer 27 Perturbed answer: Step 1: Identify theSharper Information In this problem, we are given theSharper Information about theSharper Information about Alyssa, Mike, and Tom’s limes and plums. We need to find theSharper Information about theSharper Information about theSharper …
Figure 11: Changing on the logic shifting
Statistical analysis
[…]
We first evaluate the tokens ranked by the PageRank algorithm. Specifically, we calculate the average InFidelity score when muting the corresponding neuron nodes and identify the maximum-impact neuron nodes within the top-$k$ positions. The results, summarized in Table 1, include the top-1–10 and top-11–20 nodes. Overall, the top-10 nodes show a markedly higher influence on the model’s reasoning process, indicating that the PageRank algorithm effectively identifies influential nodes in the GraphGhost framework. Moreover, larger models demonstrate greater stability than smaller ones—for instance, Llama-8B achieves an InFidelity score of 0.28, whereas Qwen3 reaches 0.58. Interestingly, the <space> token emerges as the most impactful token, exerting a strong influence on overall model performance, particularly in higher layers.
@@
We further analyze the relationship between the in-degree ratio across layers and the final reasoning predictions in the Qwen3-0.6B model, as illustrated in Figure 12. Neuron nodes with higher in-degree values are likely to merge semantic information from multiple sources. To verify this, we compute the language-changing ratio when muting individual neuron nodes across layers, defined as $R=\frac{\#\mathrm{non-English\ answer}}{\#\mathrm{All\ answer}}$.where a non-English answer refers to any generated response that switches from English to another language after perturbation. As shown in Figure 12(a), the peaks of the language-changing ratio align closely with layers exhibiting high in-degree ratios. These layers are primarily distributed in the upper regions of the model, which also correspond to the increase in InFidelity scores observed in Figure 12(b). This pattern suggests that neuron nodes in higher layers are more likely to connect directly to logit neurons, thereby altering the logical flow of reasoning (e.g., causing generation loops) rather than simply substituting equivalent semantic representations.
@@ 5 Conclusion
In this work, we introduced GraphGhost, a framework that models neuron activations and their propagation as graphs, providing a structural view of LLM reasoning. Using graph algorithms and structural interventions, we show how key neurons shape the logical flow and semantic consistency. These findings establish GraphGhost as an interpretation framework for understanding the structural foundations of reasoning in LLMs.
@@
Appendix A Related work
A.1 Interpretation
Interpreting how LLMs perform various reasoning tasks is an important step toward aligning them with human understanding Yao et al. (2023b); Zhao et al. (2024). Current work mainly falls into two directions: self-explanation Gu et al. (2024); Koa et al. (2024) and uncovering the underlying causal mechanisms Bloom et al. (2024); Dunefsky et al. (2024). The self-explanation approach, such as Chain-of-Thought prompting Wei et al. (2022); Kojima et al. (2022), requires LLMs to verbalize their reasoning steps, which can help guide them to correct answers. However, there is ongoing debate about whether such reasoning chains are truly faithful Turpin et al. (2023); Barez et al. (2025).
A.2 Reasoning tasks
Reasoning tasks are among the most important benchmarks for evaluating the ability of LLMs to solve complex problems. Specifically, they assess model performance across mathematical, scientific, and logical reasoning. Mathematical reasoning evaluates a model’s capacity to perform structured, multi-step numerical computations expressed in natural language Cobbe et al. (2021); Hendrycks et al. (2021). Scientific reasoning tasks require LLMs to draw on background knowledge and apply it in reasoning processes Khot et al. (2020); Clark et al. (2018). Logical reasoning tasks assess whether models can apply formal rules rather than relying on surface-level correlations Clark et al. (2019); Saparov and He (2022). Recent research further suggests that LLMs are capable of transferring reasoning strategies across diverse tasks, indicating that despite differences in domain, the underlying reasoning logic may be shared.
Appendix B Algorithm
Here we present the algorithm of construct the GraphGhost in Algorithm 1.
Algorithm 1 GraphGhost Construction
1:Input: A set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each $s_{i}$ contains an answer token $t_{n}^{(i)}$
2:Initialize global GraphGhost $G=(V,E,W)$, $V=\emptyset$, $E=\emptyset$, $W=\emptyset$
3:for each sample $s_{i}\in\mathcal{S}$ do
4:  Initialize token list $T=\{t_{n}^{(i)}\}$
@@ Appendix B Algorithm
6:   if Token $t$ not in question of $s_{i}$ then
7:     # trace contributing tokens and local attribution graph
8:     $\hat{T},G_{A}^{\prime}\leftarrow\text{CircuitTracer}(t)$
@@
10:     # update global GraphGhost with local graph
11:     $\hat{V}=\{v^{1}(t_{1}),\dots,v^{\ell}(t_{n})\},\hat{E}\leftarrow G_{A}^{\prime}$
12:     $V\leftarrow V\cup\hat{V}$
13:     for each edge $e$ in $\hat{E}$ do
14:      if $e\notin E$ then
15:        $E\leftarrow E\cup\{e\}$
16:        $W(e)\leftarrow 1$
17:      else
18:        $W(e)\leftarrow W(e)+1$
19:      end if
20:     end for
21:   end if
22:  end for
23:end for
24:return $G=(V,E,W)$
Appendix C Appendix: Experiment setting
C.1 Synthetic Experiment
We first generate directed graphs with 10 nodes and an edge density of 0.6, then randomly drop 60% of the edges. From each resulting subgraph, we randomly select two nodes and retain the subgraph if the path length between them is 3. In total, we collect 100,000 such graph cases. For training, we use a 5-layer decoder-only transformer with a hidden size of 96 and 12 attention heads. The model is trained for 20,000 iterations to ensure stable convergence.

### v2

@@ 3.2.1 Method
3.2.2 Analysis
Beyond visualization, we conduct further graph-based analyses to examine how the structures identified by GraphGhost correspond to LLM behaviors. Specifically, we analyze the structures from two perspectives: (1) token-merging behavior, quantified by node in-degree; and (2) trigger tokens that initiate reasoning, identified using PageRank. In addition, we present case studies to illustrate how perturbations based on GraphGhost-identified structures affect the progression of the reasoning process.
We consider Qwen3-0.6B (28 layers), Qwen2.5-1.5B (28 layers), Llama3.1-8B-Instruct (32 layers), DeepSeek-distilled-Qwen2.5B-1.5B (DS-Qwen-1.5B; 28 layers), and DeepSeek-distilled-Llama-8B (DS-Llama-8B; 32 layers). We evaluate these models on three reasoning domains: (i) mathematical reasoning (GSM8K, MAWPS), (ii) logical reasoning (ProntoQA, BoolQA), and (iii) scientific reasoning (ARC-Easy, QASC). We obtain the GraphGhosts for each model and dataset. All experiments are conducted on a single H200 GPU. Each sample will run 0.5-3h, depended on the length of reasoning tokens. Full experimental details are provided in Appendix B.2.
Semantic Merging
We begin by analyzing semantic merging behavior using the average in-degree of neuron nodes across layers. We assume that multiple tokens converge within the network, indicating potential semantic consolidation. For instance, in the representation of the number “10”, the tokens “1” and “0” exhibit convergent interactions in intermediate layers, consistent with the formation of a unified numerical representation.
In the GraphGhost representation, in-degree quantifies how many contributing nodes are merged into a target node. Accordingly, nodes with higher in-degree correspond to tokens that integrate information from multiple sources and serve as central aggregation points for reasoning-related signals. Based on this, we examine which layers and tokens are preferentially in LLMs by GraphGhost structures.
The statistical results are summarized in Appendix C and visualized in Figure 6. We observe that tokens with high in-degree exhibit clear domain sensitivity. For example, in the ProntoQA dataset, the model is more sensitive to the period token than to the comma, suggesting that punctuation plays different semantic roles across reasoning domains. Across models, however, the dominant tokens are largely consistent: special tokens such as <space>, the period, and “the” consistently rank among the top, and this pattern persists even after finetuning.
@@ 3.2.2 Analysis
Figure 6: The top merging nodes evaluated by in-degree ratio. Left: The highest score of the in-degree ratio across datasets. Right: The highest score of the in-degree ratio across the models
We further set the in-degree ratio to measure the merging behavior across the datasets and models. Let $d_{i}^{(\ell)}$ denote the in-degree of neuron node $(t_{i},\ell)$, where $\ell$ is the layer index and $i$ indexes the neuron within that layer. The in-degree ratio of layer $\ell$ is then defined as $R_{\ell}=\frac{\sum_{i=1}^{N_{\ell}}d_{i}^{(\ell)}}{\sum_{k=1}^{L}\sum_{i=1}^{N_{k}}d_{i}^{(k)}},$ where $N_{\ell}$ is the number of neurons in layer $\ell$ and $L$ is the total number of layers. A higher in-degree ratio $R_{\ell}$ implies that layer $\ell$ concentrates a larger fraction of merging behavior, indicating that many reasoning-related signals converge at this layer. In contrast, a lower ratio suggests that neurons in the layer are less involved in integrating information, playing a more distributive or transitional role in the reasoning process.
@@
As shown in Figure 7, the in-degree ratio distribution is imbalanced across both models and tokens. Although Qwen models and the Llama model perform semantic merging primarily in the bottom and top layers, the exact layers differ between them. Similarly, token merging patterns are not always consistent. Most tokens are merged in shallow layers, whereas high in-degree tokens such as period and “the" tend to merge at higher layers.
@@
Figure 8: Top contributing tokens and layers in the dataset-level GraphGhost of Qwen3-0.6B across different datasets. Left: The most frequently contributing tokens. Right: The most frequently contributing tokens.
Figure 9: Top contributing tokens and layers across models evaluated by MAWPS dataset. Left: The most frequently contributing tokens. Right: The most frequently contributing tokens.
Case studies Semantic merging controls language Question: A ) nervous and digestive B ) nervous and circulatory C ) respiratory and digestive D ) respiratory and circulatory Question: Which two body systems would most directly remove extra fluid from a person’s lungs? Original answer: … B) Nervous and circulatory systems: The nervous system is responsible for regulating the body’s response to stress, while the circulatory system is responsible for transporting blood throughout the body. Neither of these systems directly remove extra fluid from the lungs… Mute the neuron node ’the’ at layer 20 Perturbed answer: … B) Nervous and circulatory systems: The nervous system is responsible for controlling the - circulatory system is responsible for transporting blood throughout the分区肺. While the分区肺需要血液来清除多余的水分,这些系统本身并不直接参与这个过程。…   Reasoning trigger changes logic Question: There are 7.0 crayons in the drawer and 6.0 crayons on the desk. Sam placed 4.0 crayons and 8.0 scissors on the desk. How many crayons are now there in total? Let’s think step by step (ground truth: 17) Original answer: … - Sam placed 4.0 crayons on the desk. - Sam also placed 8.0 scissors on the desk. …. - Total crayons on the desk = 7.0 + 4.0 + 8.0 - Total crayons on the desk = 19.0 Mute the neuron node ’.’ at layer 25 Perturbed answer: … - Sam placed 4.0 crayons on the desk. - So, the total number of crayons on the desk is now 6.0 (initial crayons) + 4.0 (Sam’s crayons) = 10.0 crayons. … - So, the total number of crayons = 7.0 (drawer) + 10.0 (desk) = 17.0 crayons.
Figure 10: Semantic control and reasoning triggers change the model generation processing. In the language-switching case, the generated characters are direct translations of the original answer.
Reasoning Triggers
The reasoning structure is segmented by special tokens such as So, Step, and Wait Bogdan et al. (2025). We assume that activated neuron nodes in the middle layers trigger these special token expressions at the logit layers. At the same time, the neuron nodes themselves are activated by input tokens. To capture both effects, we treat logit-layer nodes as signals directed back to the neuron nodes by reversing their edge directions. We then apply the PageRank algorithm to identify the most influential nodes. We summarize the results in Figure 8 and Figure 9, and the full details can be found in Appendix C.
@@
Case Study
We further conduct case studies to illustrate how semantic merging and reasoning-trigger effects influence the reasoning process. Based on the analyses in the previous sections, we observe that tokens, such as “the” and “.”, are identified as important by both analysis methods, although their functional roles vary across layers. For instance, in Qwen3, these tokens tend to merge in middle layers (e.g., around layer 15), while in higher layers (e.g., beyond layer 25), they often act as reasoning triggers. Accordingly, we design case studies to examine how the roles of such tokens evolve across layers during reasoning.
Specifically, we mute neurons in the LLM when they correspond to a given token at a specified layer. We show that muting neurons in middle layers leads to semantic changes, as illustrated in Figure 10. When the neuron node associated with “the” is deleted at layer 20, the English semantic pathway is disrupted, activating Chinese neurons that share similar semantic signals. As a result, the model produces output in Chinese, effectively translating from the original English text. In the Reasoning trigger case, deleting the neuron node “.” at layer 25 introduces a new reasoning token such as “So", ultimately leading the model to produce the correct answer. The full results are shown in Appendix B.3.
@@
4 Quantitative Experiments
To evaluate whether the structural patterns identified within LLMs are functionally relevant to reasoning behavior, We perform quantitative experiments that examine whether GraphGhost-identified tokens exhibit consistent and measurable significance. We first introduce the evaluation metrics in Section 4.1. The evaluation considers both the sample-level view (Section 4.2) and the dataset-level view (Section 4.3).
4.1 Metrics
[…]
4.2 Contribution Token Selection in GraphGhost: Sample View
To evaluate whether GraphGhost can identify tokens that are functionally important for the reasoning process, we compare it against attention-based explanation methods (Clark et al., 2019; Liu et al., 2024) and random baselines. Instead of analyzing individual tokens, we adapt the same pipeline as GraphGhost to evaluate its effectiveness. The detailed settings are provided in Appendix B.4.
The comparative results are summarized in Figure 11. For a detailed breakdown of the performance for each individual model without averaging, we provide the complete results in Appendix D (Figure 15). GraphGhost consistently outperforms both random and attention-based baselines. This advantage is most pronounced on mathematical reasoning tasks (MAWPS and GSM8K), where the attention baseline performs comparably to random noise, indicating its failure to capture essential reasoning structures. In contrast, GraphGhost maintains high fidelity even under low sparsity constraints. Moreover, on the challenging ARC dataset, GraphGhost demonstrates superior robustness; notably, it is the only method that surpasses the 0.5 fidelity threshold on Qwen3. These results suggest that while attention primarily captures surface-level correlations, GraphGhost is able to trace deeper, multi-hop structural dependencies that are critical for complex reasoning.
4.3 Behaviors Controlled by Critical Nodes in GraphGhost: Dataset View
@@
We first analyze the relationship between the in-degree ratio across layers and the final reasoning predictions in the Qwen3-0.6B model, as illustrated in Figure 12. Neuron nodes with higher in-degree values are likely to merge semantic information from multiple sources. To verify this, we compute the language-changing ratio when muting individual neuron nodes across layers, defined as $R=\frac{\#\mathrm{non-English\ answer}}{\#\mathrm{All\ answer}}$, where a non-English answer refers to any generated response that switches from English to another language after perturbation. As shown in Figure 12(a), the peaks of the language-changing ratio align closely with layers exhibiting high in-degree ratios in middle layers. Furthermore, as shown in Figure 12(b), Infidelity increases in the upper layers. This indicates that neuron nodes in higher layers are more likely to connect directly to logit neurons, thereby altering the underlying reasoning logic. This observation is consistent with the findings in Section 3.2.2, where perturbations in middle layers primarily affect semantic expression, whereas perturbations in higher layers tend to modify the reasoning logic itself.
We also evaluate the important tokens ranked by PageRank algorithm. We calculate the average InFidelity score when muting the corresponding neuron nodes and identify the maximum-impact neuron nodes within the top-$k$ positions. The results, summarized in Table 1, include the top-1–10 and top-11–20 nodes. Overall, the top-10 nodes show much higher influence, indicating that PageRank effectively identifies important nodes in GraphGhost. Larger models are more stable than smaller ones; Llama-8B achieves an InFidelity score of 0.28, whereas Qwen3 reaches 0.58. Notably, the <space> token is the most influential, especially in higher layers.
@@ 5 Conclusion
We present GraphGhost, a graph-based framework that reveals the internal structures underlying LLM reasoning. By linking token interactions and neuron activations to graph properties, GraphGhost shows how semantic merging and reasoning triggers shape model behavior. Our results demonstrate that LLM reasoning is governed by structured internal graphs, offering a principled path toward mechanistic interpretability.
@@
Potential Risk
A potential risk of our method is that the extracted graph structures may be over-interpreted as definitive explanations of model behavior, while they only provide approximate views of internal information flow. We caution against such misuse and position our approach as a complementary analytical tool.
Appendix A Appendix: Algorithm
Here we present the algorithm of construct the sample and dataset view of GraphGhost in Algorithm 1 and Algorithm 2.
Algorithm 1 GraphGhost Construction (Sample View)
1:Input: A sample $s$ containing a final answer token $t_{n}$
2:Output:
  • Sample-level GraphGhost $G^{\text{sample}}=(V,E)$ (unweighted)
  • A local edge-weight map $W^{\text{local}}$ (for dataset aggregation only)
3:Initialize $V=\emptyset$, $E=\emptyset$, and $W^{\text{local}}(e)=0$ for all edges $e$
4:Initialize frontier $T=\{t_{n}\}$
@@ Appendix A Appendix: Algorithm
6:  if $t$ is not part of the question tokens then
7:   # Step 1: Local logit-based attribution graph
8:   $(\hat{T},G_{A}^{\prime})\leftarrow\textsc{CircuitTracer}(t)$ $\triangleright$ $G_{A}^{\prime}=(\hat{V},\hat{E},W^{\text{loc}})$
@@
10:   # Step 2: Add edges to sample-level graph (structure only)
11:   $V\leftarrow V\cup\hat{V}$
12:   $E\leftarrow E\cup\hat{E} \triangleright$ unweighted sample graph
13:   # Step 3: Accumulate local weights for dataset view
14:   for each edge $e\in\hat{E}$ do
15:     $W^{\text{local}}(e)\leftarrow W^{\text{local}}(e)+W^{\text{loc}}(e)$
16:   end for
17:  end if
18:end for
19:return $(G^{\text{sample}}=(V,E),\;W^{\text{local}})$
Algorithm 2 GraphGhost Construction (Dataset View)
1:Input: A set of samples $\mathcal{S}=\{s_{1},\dots,s_{m}\}$
2:Output: Dataset-level GraphGhost $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$
3:Initialize $V^{\text{data}}=\emptyset$, $E^{\text{data}}=\emptyset$, and $W^{\text{data}}(e)=0$
4:for each sample $s_{i}\in\mathcal{S}$ do
5:  $(G^{(i)},W^{\text{local}})\leftarrow\textsc{GraphGhost-SampleView}(s_{i})$
6:  $(V^{(i)},E^{(i)})\leftarrow G^{(i)}$
7:  # Merge sample graph structure
8:  $V^{\text{data}}\leftarrow V^{\text{data}}\cup V^{(i)}$
9:  $E^{\text{data}}\leftarrow E^{\text{data}}\cup E^{(i)}$
10:  # Accumulate edge weights across samples
11:  for each edge $e\in E^{(i)}$ do
12:   $W^{\text{data}}(e)\leftarrow W^{\text{data}}(e)+W^{\text{local}}(e)$
13:  end for
14:end for
15:return $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$
Appendix B Appendix: Experiment setting
B.1 Synthetic Experiment
We first generate directed graphs with 10 nodes and an edge density of 0.6, then randomly drop 60% of the edges. From each resulting subgraph, we randomly select two nodes and retain the subgraph if the path length between them is 3. In total, we collect 100,000 such graph cases. For training, we use a 5-layer decoder-only transformer with a hidden size of 96 and 12 attention heads. The model is trained for 20,000 iterations to ensure stable convergence where the validation loss is the minimal.